# Reinforcement Learning with Intrinsically Motivated Feedback Graph for Lost-sales Inventory Control

### v1

@@ 3.3 Theoretical Analysis
We conduct a quantitative analysis of how the sample complexity is reduced with the FG in the lost-sales IC environment. The qualitative analysis based on the property of RLFG is in Appendix 7. Here we provide an analysis based on Q-learning for simplicity. The detailed proof is in Appendix 8
Without loss of generality, we restrict some definitions in section 2.3 and define some new concepts. We consider the reward function $R:\mathcal{S}\times\mathcal{A}\to(0,1)$. The demand $d_{t}\sim P_{d}(d)$ can obey any independent discrete distribution and $d^{\text{max}}<y^{\text{max}}$. We define $\pi_{b}$ as the stationary behavior policy and $\mu(\bm{s},a)$ as the stationary distribution of the Markov chain under $\pi_{b}$ and $P(\bm{s}^{\prime}|\bm{s},a)$, which is the same as the update probability in typical RL. In RLFG, the stationary distribution does not change but the update probability becomes $\tilde{\mu}(\bm{s},a)\geq\mu(\bm{s},a)$ because of the side experiences.
Scenario 1: Consider a graph $\mathcal{G}$ with all state-action pairs as the nodes and there is no edge between the nodes. It can be regarded as $\mathcal{G}=\mathcal{G}_{1}\cup\mathcal{G}_{2}$, where $\mathcal{G}_{1}=\emptyset$ and $\mathcal{G}_{2}$ consists of nodes without any edge.
Lemma 1: The sample complexity of the asynchronous Q-learning under scenario 1 is analyzed in Li et al. (2020), which is $\tilde{O}(\frac{1}{\mu_{\text{min}}(1-\gamma)^{5}\epsilon^{2}}+\frac{t_{\text{mix}}}{\mu_{\text{min}}(1-\gamma)})$, where $\mu_{\text{min}}=\min_{(s,a)\in\mathcal{G}}\mu(\bm{s},a)$ and $t_{\text{mix}}$ is the mixing time of the chain.
Lemma 1 indicates that the sample complexity of Q-learning without FG is determined by $\mu_{\text{min}}$. We will show how Q-learning with FG improves $\mu(\bm{s},a)$ to reduce the sample complexity.
Scenario 2: Consider a graph $\mathcal{G}$ with all state-action pairs as the nodes. Assume for nodes satisfying $y\geq d_{t}$, once one node is sampled, all of these nodes can be sampled and updated simultaneously. For nodes with $y<d_{t}$, only nodes with $y^{\prime}\leq y$ can be sampled and updated simultaneously. Thus $\mathcal{G}$ can be regarded as $\mathcal{G}=\mathcal{G}_{1}\cup\mathcal{G}_{2}$.
Theorem 1: The update probability for Q-learning with FG under scenario 2 is:
$\tilde{\mu}(\bm{s},a)=\underbrace{\sum_{d=0}^{d^{\text{max}}}P_{d}(d)\sum_{\begin{subarray}{c}(\bm{\bar{s}},\bar{a})\in\mathcal{G}\\ \bar{y}\geq d\end{subarray}}\mu(\bm{\bar{s}},\bar{a}|d)}_{\text{Uncensored term}}+\underbrace{\sum_{d=0}^{y-1}P_{d}(d)\sum_{\begin{subarray}{c}(\bm{\bar{s}},\bar{a})\in\mathcal{G}\\ y\leq\bar{y}\leq d\end{subarray}}\mu(\bm{\bar{s}},\bar{a}|d)}_{\text{Censored term}}.$ (5)
Conclusion: We can obtain the relationship between $\tilde{\mu}(\bm{s},a)$ and $\mu(\bm{s},a)$ in Equation (6). It shows that FG not only improves $\mu_{\text{min}}$ but also improves that of all the state-action pairs.
@@ 3.4 Intrinsically Motivated Exploration
To further utilize the benefits of FG, we design an intrinsic reward by incorporating the information of the side experiences into the curiosity-driven exploration. In particular, for a state-action pair, the number and the average uncertainty of the side experiences generated by this state-action pair are incorporated into the intrinsic reward. With this intrinsic reward, the agent is directed towards the state-action space where more side experiences can be generated.
This idea is inspired by the theoretical analysis based on Equation 5. The uncensored term indicates how much the uncensored case contributes to improving the probability of being updated. This term contributes to all state-action pairs so improving this term can improve the probability of being updated for all state-action pairs. The censored term indicates how much benefit the current state-action pair can obtain from the censored cases. For the censored case, the sample complexity can be further reduced by visiting the state-action pairs with larger inventory. To satisfy both conditions, designing a behavior policy manually is difficult and not general enough. Thus we design an intrinsic reward, written in Equation 7 because both conditions lead to generating more side observations.
$r^{in}_{i}=r^{in}_{i}+log_{10}(J)\times\frac{1}{J}\sum_{j=1}^{J}r^{in}_{i,j}.$ (7)
Algorithm 3 shows the details of the intrinsic reward module. We utilize the M-head DQN (Nikolov et al. 2018) with each head trained by different mini-batch of experiences to get the prediction error as curiosity rewards. The final intrinsic reward, denoted in Equation 7, consists of the curiosity reward of the experience itself ($r^{in}_{i}$) and the averaged curiosity reward of all corresponding side experiences ($r^{in}_{i,j}$) generated by this experience. The averaged value preserves the advantage of scale invariance but loses the quantity information compared with the sum value. To balance these two aspects, we multiply the averaged value with $\log_{10}(J)$ to incorporate more quantity information into the intrinsic reward. In the censored condition, where $J=y_{t}$, a larger $J$ increases the intrinsic reward, which means more side experiences. In the uncensored condition, $J=y^{\text{max}}$ is larger than $y_{t}$, which means that the intrinsic reward is more likely to be larger than that in the censored condition.
@@
4 Experiment
To verify the performance and sample efficiency of our decision framework, we utilize the example algorithm in Algorithm 1 and test it based on the standard lost-sales IC environment. All experiments are averaged with 20 random seeds with shaded areas representing the standard deviation.
4.1 Setup
4.1.1 Baseline
Heuristic Methods: We compare our method with widely recognized and well-performed heuristic methods. All parameters used in these methods are searched in detail and the best results are selected.
  • Constant Order: The order is always a constant ($a_{t}=r^{h}$), where $r^{h}$ is a parameter.
  • Myopic 1-period (Morton 1971): Assume the distribution of $d_{t}$ is known. This method aims to minimize the expected cost at $t+L$. The order is $a_{t}=\min_{P(u_{t+L}<0)\leq\frac{c+h}{p+h}}a_{t}$, $0\leq a_{t}\leq a^{\text{max}}$.
  • Myopic 2-period (Zipkin 2008a): Since Myopic 1-period only considers the expected cost for one time step, it can be improved by considering 2-time steps.
  • Base-Stock method (Zipkin 2008a): The Base-Stock method aims to keep the inventory level constant, including upcoming orders. The order is $a_{t}=(S^{h}-\bm{1}\cdot\bm{s}_{t})^{+}$, where $S^{h}$ is a parameter.
  • Capped Base-Stock method (Xin 2021): It combines Base-Stock and Constant Order method. The order is $a_{t}=\min[(S^{h}-\bm{1}\cdot\bm{s}_{t})^{+},r^{h}]$, where $S^{h}$ and $r^{h}$ are parameters.
  • Bracket method (Bai et al. 2023): A variant of the Constant Order method. The order is $a_{t}=\lfloor(t+1)r^{h}+\theta^{h}\rfloor-\lceil tr^{h}+\theta^{h}\rceil$, where $r^{h}$ and $\theta^{h}$ are parameters.
Deep Reinforcement Learning Methods: We choose to compare Rainbow-FG with the A3C method in Gijsbrechts et al. (2022), which is an on-policy RL method; Rainbow, which is the basis of Rainbow-FG; Rainbow-FG(H), which utilizes the heuristic knowledge to continue finetuning Rainbow-FG. Rainbow-FG(H) can be regarded as the upper bound of Rainbow-FG. Since A3C is on-policy, it is not applicable to compare sample efficiency with off-policy Rainbow-FG. Thus A3C is only used to compare the optimal results.
Table 2: Parameters of the environment. $L$ and $p$ are two important parameters where $L$ reflects the dimensionality of the state and $p$ reflects the seriousness of lost sales.
Parameter $L$ $p$ $y^{\text{max}}$ $a^{\text{max}}$ $c$ $h$ $d^{m}$ $d^{\text{max}}$
Value $4$ $4$ $100$ $20$ 00 $1$ $5$ $20$
4.1.2 Environments
We test our method and baselines across different settings of the standard discrete lost-sales IC environment with demand obeying Poisson distribution, where $d^{m}$ indicates the mean value. The parameters of the environment setting are shown in Table 2 and the hyperparameters of our method are in Appendix 9. As for parameters in Table 2, we follow the common recognized settings according to Zipkin (2008a), Xin (2021), and Gijsbrechts et al. (2022). In the following sections, default parameters are used unless stated otherwise. The hyperparameter analysis is demonstrated in Appendix 12.
Table 3: Optimal result comparison. For each method, the first row is the optimal average cost and the second row is the optimality gap. The bold values indicates the values of the best four methods.
@@ 4.1.2 Environments
4.2 Optimal Results Comparison
@@
Table 3 presents the average results and optimality gap compared to the optimal results. Without strong heuristics, Rainbow demonstrates similar performance to the A3C method in Gijsbrechts et al. (2022), whereas Rainbow-FG achieves superior results. This notable improvement can be attributed to Rainbow-FG’s ability to leverage a broader range of side experiences from the feedback graph, enabling it to converge towards better solutions compared with Rainbow.
@@
When contrasting Rainbow-FG with the heuristic methods, Rainbow-FG still falls short of surpassing the top-performing heuristic methods, such as Myopic 2-period and Capped Base-Stock. However, Myopic 2-period method assumes the distribution of the demand is known, which is a strong assumption, and the Capped Base-Stock method needs extensive parameter searches to attain this optimal result. These parameters vary across different scenarios without patterns. Conversely, Rainbow-FG consistently achieves close performance and possesses adaptability to diverse settings without these strong heuristics. Furthermore, after incorporating the heuristic knowledge, Rainbow-FG(H) attains equivalent or even superior performance compared to the best heuristic methods, particularly in certain settings. This result indicates that Rainbow-FG has the potential to perform better than heuristic methods.
((a)) $\epsilon=0$
((b)) $\epsilon=0.1$
((c)) $\epsilon=0.4$
Figure 2: Learning process of Rainbow and Rainbow-FG under the IC environment. “Final” denotes the optimal result of Rainbow-FG. The first row is the initial-stage view and the second row is the near-convergence view of the learning process.
4.3 Sample Efficiency Comparison for FG
@@
Figure 2 illustrates the learning process of Rainbow and Rainbow-FG with different exploration parameters. The utilization of FG significantly enhances the sample efficiency. During the initial stage, Rainbow-FG with all $\epsilon$ consistently outperforms Rainbow. Towards the end, Rainbow-FG with $\epsilon=0$ converges to the final result first and subsequently trails Rainbow-FG with $\epsilon=0.1$. Conversely, Rainbow fails to converge to the final result throughout the entire 100 episodes. Furthermore, FG contributes to improving the stability of the learning process. Notably, the learning process of Rainbow-FG exhibits lower standard deviation and more stable learning curves. Under low exploration conditions ($\epsilon=0$), Rainbow-FG demonstrates significantly faster learning compared to Rainbow. This outcome signifies the benefits of FG when exploration is not favorable. Experiments on more settings are shown in Appendix 11.
4.4 Sample Efficiency Comparison for Intrinsic Reward
This section investigates the effect of intrinsic reward designed for the IC problem. Figure 3 illustrates the learning process of Rainbow-FG with and without the intrinsic reward, referred to as Rainbow-FG w/ inr and Rainbow-FG w/o inr, respectively. The results demonstrate that the designed intrinsic reward significantly improves sample efficiency during the initial stages of training in the environment with $p=4,L=4$, and $d^{m}=5$. However, as the learning process reaches the Constant Order level, the intrinsic reward does not exhibit a substantial impact. We attribute it to the simplicity of the environment, which weakens the effect of the intrinsic reward. To further evaluate its effectiveness, we conduct experiments in more complex settings with $p=19,L=8$, and varying values of $d^{m}=$ (5, 10, 15). In these environments, we observe a clear effect of the intrinsic reward during the whole training process. Moreover, Rainbow-FG w/ inr demonstrates a more stable training process compared to Rainbow-FG w/o inr, achieving slightly lower cost.
((a)) p=4, L=4, $d^{m}$=5
((b)) p=19, L=8, $d^{m}$=5
((c)) p=19, L=8, $d^{m}$=10
((d)) p=19, L=8, $d^{m}$=15
Figure 3: Learning process of Rainbow-FG with and without intrinsic reward. The two rows have the same meaning with Figure 2.

### v2

@@ 3.3 Theoretical Analysis
We conduct a quantitative analysis of how FG reduces sample complexity in the lost-sales IC environment. Here we provide an analysis based on Q-learning for simplicity. The qualitative analysis based on the property of RLFG is in Appendix 8.
Without loss of generality, we restrict some definitions in Section 2.2 and define some new notations. We consider the reward function $R:\mathcal{S}\times\mathcal{A}\to(0,1)$. The demand $d_{t}\sim P_{d}(d)$ follows any independent discrete distribution and $d^{\text{max}}<y^{\text{max}}$. We define $\pi_{b}$ as the stationary behavior policy and $\mu(\bm{s},a)$ as the stationary distribution of the Markov chain under $\pi_{b}$ and $P(\bm{s}^{\prime}|\bm{s},a)$. In standard RL, $\mu(s,a)$ also indicates the update probability since only visited state-action pairs can be updated.
Scenario 1: Consider an edgeless graph $\mathcal{G}$ with all state-action pairs as the nodes.
Lemma 1: The sample complexity of the asynchronous Q-learning under scenario 1 is analyzed by Li et al. (2020), which is $\tilde{O}(\frac{1}{\mu_{\text{min}}(1-\gamma)^{5}\epsilon^{2}}+\frac{t_{\text{mix}}}{\mu_{\text{min}}(1-\gamma)})$, where $\mu_{\text{min}}=\min_{(s,a)\in\mathcal{G}}\mu(\bm{s},a)$ and $t_{\text{mix}}$ is the mixing time of the Markov chain.
Lemma 1 indicates that the sample complexity of Q-learning without FG is determined by $\mu_{\text{min}}$. In the following parts, we will demonstrate: Although the stationary distribution is unchanged, RLFG improves sample efficiency because of increased update probability ($\tilde{\mu}(\bm{s},a)\geq\mu(\bm{s},a)$) due to the side experiences. From scenario 1 to scenario 6, we generalize the condition from RL to RLFG, which is provided in Appendix 9. Here we directly show the final condition of RLFG in scenario 6.
Scenario 6: Consider a graph $\mathcal{G}$ with all state-action pairs as the nodes. Consider all nodes satisfying $y\geq d_{t}$, if one is sampled by the policy, all can be sampled by the FG module and updated simultaneously. For nodes with $y<d_{t}$, only nodes with $y^{\prime}\leq y$ can be sampled by the FG module and updated simultaneously.
Theorem 1: The update probability for Q-learning with FG under scenario 6 is Equation (5). The proof is provided in Appendix 9.
$\tilde{\mu}(\bm{s},a)$ $=\underbrace{\sum_{d=0}^{d^{\text{max}}}P_{d}(d)\sum_{\begin{subarray}{c}(\bm{\hat{s}},\hat{a})\in\mathcal{G}\\ \hat{y}\geq d\end{subarray}}\mu(\bm{\hat{s}},\hat{a}|d)}_{\text{Uncensored term}}+\underbrace{\sum_{d=0}^{y-1}P_{d}(d)\sum_{\begin{subarray}{c}(\bm{\hat{s}},\hat{a})\in\mathcal{G}\\ y\leq\hat{y}\leq d\end{subarray}}\mu(\bm{\hat{s}},\hat{a}|d)}_{\text{Censored term}}.$ (5)
Remark: We can obtain the relationship between $\tilde{\mu}(\bm{s},a)$ and $\mu(\bm{s},a)$ in Equation (6), which shows that FG improves the update probability for all the state-action pairs and thus improves the $\mu_{\text{min}}$, which is the state-action pair with minimum of update probability.
@@ 3.4 Intrinsically Motivated Exploration
To further leverage the benefits of FG, we design a curiosity-driven and FG-driven intrinsic reward to direct the RL algorithm towards the state-action space where more side experiences can be generated. For each state-action pair, the quantity and average uncertainty of the side experiences generated by this state-action pair are incorporated into the intrinsic reward.
This idea is motivated by the theoretical analysis based on Equation 5. The uncensored term reflects the contribution of the uncensored case to improve the update probability. Enhancing this term can improve the update probability of all state-action pairs. The censored term measures the benefits a state-action pair gains from the censored cases. In this scenario, the update probability can be improved from the state-action pairs with larger inventory levels. To satisfy both conditions, manually designing a behavior policy is difficult and not general enough. Instead, we propose an intrinsic reward given in Equation 7 promoting side experience generation.
$r^{in}_{i}=r^{in}_{i}+log_{10}(J)\times\frac{1}{J}\sum_{j=1}^{J}r^{in}_{i,j},$ (7)
where $J$ is the number of the side experiences generated by an experience. Algorithm 2 shows the details of the intrinsic reward module. We utilize the M-head DQN (Nikolov et al. 2018) with each head trained by different mini-batches of experiences to get the prediction error as the curiosity. The final intrinsic reward, denoted in Equation 7, consists of the curiosity of the experience itself ($r^{in}_{i}$) and the averaged curiosity of all corresponding side experiences ($r^{in}_{i,j}$). The averaged value preserves the advantage of scale invariance but loses the quantity information compared with the sum value. To balance these two aspects, we multiply $\log_{10}(J)$ with the averaged value to incorporate more quantity information into the intrinsic reward. In the censored case, where $J=y_{t}$, a larger $J$ increases the intrinsic reward, which means more side experiences. In the uncensored case, $J$ has no differences, and a larger intrinsic reward indicates higher curiosity. Across the censored and uncensored case, $J=y^{\text{max}}$ is larger than $y_{t}$, which means that the intrinsic reward in the uncensored case is more likely to be larger than that in the censored case.
@@
Table 2: Optimal result comparison on single-item environment. For each method, the first row is the optimal average cost ($\downarrow$) and the second row is the optimality gap ($\downarrow$). The bold values indicates the values of the best five methods. $p$ denotes the lost-sales penalty, and $L$ denotes the lead time.
@@
4 Experiment
To verify the performance and sample efficiency of our training framework, we test our method, including (1) Rainbow-FG, (2) TD3-FG, and (3) Rainbow-FG (H), on the (a) single-item, (b) multi-item, and (c) multi-echelon lost-sales IC environment. Rainbow-FG (H) is a variant, which utilizes the heuristic knowledge to continue finetuning Rainbow-FG. It can be regarded as the upper bound of Rainbow-FG. All experiments are averaged on 20 random seeds with shaded areas representing the standard deviation.
4.1 Baselines
Heuristic Methods: We compare our method with widely recognized and well-performed heuristic methods with optimized parameters via brute force search.
  • Constant Order: Fixed order $a_{t}=r^{h}$ with $r^{h}$ being a parameter.
  • Myopic 1-period (Morton (1971)): Assume the distribution of $d_{t}$ is known. This method aims to minimize the expected cost at $t+L$. The order is $a_{t}=\min_{P(y_{t+L}-d_{t+L}<0)\leq\frac{c+h}{p+h}}a_{t}$, where $0\leq a_{t}\leq a^{\text{max}}$.
  • Myopic 2-period (Zipkin (2008a)): This method considers 2-step expected costs from $t+L$ compared with Myopic 1-period method.
  • Base-Stock method (Zipkin (2008a)): This method aims to maintain the inventory level with considering upcoming orders. The order is $a_{t}=(S^{h}-\bm{1}\cdot\bm{s}_{t})^{+}$, where $S^{h}$ is a parameter.
  • Capped Base-Stock method (Xin (2021)): It combines Base-Stock and Constant Order method. The order is $a_{t}=\min[(S^{h}-\bm{1}\cdot\bm{s}_{t})^{+},r^{h}]$, where $S^{h}$ and $r^{h}$ are parameters.
  • Bracket method (Bai et al. (2023)): A variant of the Constant Order method. The order is $a_{t}=\lfloor(t+1)r^{h}+\theta^{h}\rfloor-\lceil tr^{h}+\theta^{h}\rceil$, where $r^{h}$ and $\theta^{h}$ are parameters.
Deep RL Methods: We choose to compare (1) Rainbow; (2) TD3; (3) the A3C method in Gijsbrechts et al. (2022), which is an on-policy RL method. Since A3C is on-policy, its sample efficiency cannot be directly compared to the off-policy methods. Thus A3C only serves for comparing optimal performance.
Figure 2: Single-item.
Figure 3: Multi-item.
Figure 4: Multi-echelon.
4.2 Case One: Single Item
We test our methods and baselines across different settings of the single-item lost-sales IC environment with demand obeying Poisson distribution. The parameters of the environment and the hyperparameters of our methods are in Appendix 10. In the following sections, default parameters are used unless stated otherwise. The hyperparameter analysis is demonstrated in Appendix 13.
4.2.1 Optimal Results Comparison
@@
Table 2 presents the average results and optimality gap compared to the optimal results. Without strong heuristics, Rainbow, TD3 and A3C demonstrates similar performance whereas Rainbow-FG and TD3-FG achieves superior results. In particular, TD3-FG outperforms Myopic 2-period and Capped Base-Stock method and achieve SOTA performance in some settings without any strong heuristics. This notable improvement can be attributed to the ability to leverage a broader range of side experiences from FG, enabling TD3-FG to converge towards better solutions compared with TD3, and even existing best methods.
@@
Although Rainbow-FG does not surpass the top-performing heuristic methods like Myopic 2-period and Capped Base-Stock, Myopic 2-period method assumes the distribution of the demand is known, which is a strong assumption, and the Capped Base-Stock method needs extensive parameter searches to attain this optimal result. These parameters vary across different scenarios without patterns. Conversely, Rainbow-FG consistently achieves near-optimal performance and possesses adaptability to diverse settings without these strong heuristics. Furthermore, when equipped with the heuristic knowledge, Rainbow-FG (H) attains equivalent or even superior performance compared to the best heuristic methods, particularly in settings ($p=4,L=2,3,4$). This result indicates that Rainbow-FG has the potential to perform better than heuristic methods.
4.2.2 Sample Efficiency Comparison
@@
Figure 2 illustrates the learning process of Rainbow versus Rainbow-FG and TD3 versus TD3-FG under two settings. The utilization of FG significantly enhances the sample efficiency. In the early stages of training, Rainbow-FG and TD3-FG show great sample efficiency compared with Rainbow and TD3. Towards the end, Rainbow-FG and TD3-FG converge faster and better compared to Rainbow and TD3, where Rainbow and TD3 fail to converge to the level of their FG version throughout the entire 100 episodes in some settings. Furthermore, FG contributes to improving the stability of the learning process. Notably, Rainbow-FG and TD3-FG exhibit lower standard deviation and more stable learning curves for the whole learning process. Experiments on more settings are shown in Appendix 12.
4.3 Case Two: Multi Item
We further conduct experiments with TD3 and TD3-FG on the more challenging multi-item lost-sales IC problem with 2, 3, and 5 items in one store. Note that Rainbow and Rainbow-FG is not included in the comparison due to its exponential output growth with items. The parameters of the environment are in Appendix 10.
4.3.1 Optimal Results Comparison
Figure 5 shows the final results in the multi-item environment. It demonstrates that TD3-FG can achieve better performance than TD3 in all cases. In particular, TD3-FG decreases the performance gap by 4.58%, 1.16%, and 2.56% for environments with 2, 3, and 5 items.
4.3.2 Sample Efficiency Comparison
Figure LABEL:exp-multi-item demonstrates similar learning processes of TD3 and TD3-FG with different numbers of items like sing-item environment. TD3-FG shows great sample efficiency and better stability throughout the training. As items increase, the efficiency gap between TD3 and TD3-FG widens, while TD3-FG maintains consistent gains, highlighting the scalability of the proposed methods.
4.4 Case Three: Multi Echelon
We compare TD3 and TD3-FG on multi-echelon lost-sales IC with 1 warehouse and 1, 2 or 4 retailers to evaluate generalization. The retailers order from the warehouse and satisfy the demand of customers, and the warehouse orders from the manufacturer and satisfies the demand of retailers. Both the warehouse and retailers need to find optimal decisions. The parameters of the environment are in Appendix 10.
4.4.1 Optimal Results Comparison
Figure 5 shows the final results in the multi-echelon environment. It demonstrates that TD3-FG can achieve lower costs than TD3 with lower variance. In particular, TD3 performs very poorly when the number of retailers is 4. This results from failed warehouse decisions, as its demand comprises retailer orders. Since the demand for the warehouse is the sum of the orders from retailers, the mean and variance of the demand for the warehouse also increase as the number of retailers increases. Due to failed warehouse decisions, the retailers also cannot find optimal solutions to satisfy the demand of customers. Thus both the warehouse and retailers are in a greatly lost-sales condition.
Figure 5: Final results on multi-item and multi-echelon environments.
4.4.2 Sample Efficiency Comparison
Figure LABEL:exp-multi-echelon demonstrates similar learning curves of TD3 and TD3-FG with different numbers of retailers like multi-item environment. TD3-FG always achieves better performance with smaller variance under the same training budgets. Besides, TD3-FG converges to near-optimal solutions within around 20 episodes but TD3 converges to worse solutions with 80 to 100 episodes when the numbers of retailers are 1 and 4.